# Schwarz Information Criterion Aided Multi-Armed Bandit for Decentralized Resource Allocation in Dynamic LoRa Networks

### v2

@@ I Introduction
With the rapid proliferation of the Internet of Things (IoT) devices in recent years, it has become increasingly common for a massive number of terminals to transmit data simultaneously [1]. This explosive growth has intensified channel contention in Low-Power Wide-Area Networks (LPWANs), particularly in Long Range (LoRa) networks, which stand out as a prominent example of LPWAN technology [2, 3]. To use limited wireless resources efficiently, each LoRa device must optimize its transmission parameters in response to its dynamically changing communication environment.
@@
One representative decentralized technique used in distributed resource allocation is the multi-armed bandit (MAB) algorithms [6, 7, 8, 9, 10, 11]. MAB is a reinforcement learning method that balances exploration and exploitation to learn the optimal choice, and it has been well applied to transmission parameter optimization. In particular, Upper Confidence Bound (UCB)1-tuned algorithm, which considers the variance of rewards for each option, has been widely used due to its enhanced exploration efficiency. However, most MAB based distributed transmission parameter selection methods, including UCB1-tuned, rely heavily on historical observations. This dependence results in delayed adaptation when the communication environment changes abruptly, as these algorithms lack mechanisms to automatically discard past learning history [12]. Consequently, performance degradation in dynamic environments becomes unavoidable.
@@
  • We propose a lightweight distributed learning based transmission parameter selection method. In our proposed method, only a small set of statistics need to be stored, and requires neither complex matrix operations nor centralized control. Consequently, it incurs minimal computational and memory overhead, making it well suited for implementation on memory-constrained IoT devices.
@@ II System Model
The LoRa system considered in this study is illustrated in Fig. 1, consisting of a single gateway (GW), multiple EDs, and an external interference source (jammer node). In the LoRa system, multiple channels are available in the network, and each ED selects one of them to transmit data to the GW. However, when the jammer occupies or blocks a specific channel, transmissions over that channel are more likely to experience interference. Therefore, we assume a dynamic communication environment where the jammer intermittently affects part of the frequency spectrum, causing certain channels to transition from an available state to an unavailable state during operation. Each ED must adaptively select the transmission parameters that yield the highest energy efficiency according to the varying channel conditions.
Let $G$ be the number of LoRa EDs, $M$ be the number of available channels, $P$ be the number of transmission power (TP) levels, and $L$ be the number of bandwidth (BW) options. The set of available channels is denoted as $C=\{c_{1},c_{2},\ldots,c_{M}\}$, the set of TP levels as $U=\{u_{1},u_{2},\ldots,u_{P}\}$, and the set of BW options as $B=\{b_{1},b_{2},\ldots,b_{L}\}$. Let $K=\{k_{1},k_{2},\ldots,k_{M\times P\times L}\}$ be the set of all possible combinations of channel, TP, and BW. Each ED selects one parameter set from $K$ before transmitting data based on a reinforcement learning algorithm.
@@
Furthermore, each ED records the presence or absence of ACKs for each transmission and accumulates this as a transmission history. In this study, we introduce a mechanism that statistically detects changes in the communication environment based on this history using the SIC. When a change is detected, the past learning history is discarded, and the learning process is reinitialized to enable quick adaptation to the new environment. To represent the transmission outcomes quantitatively, each ED maintains a binary observation sequence of ACK receptions as follows:
$\Omega=\{\,s_{i}\mid s_{i}\in\{0,1\},\,i=1,2,\ldots,l\,\},$ (1)
where $s_{i}=1$ denotes a successful ACK reception, and $s_{i}=0$ denotes a failure. To analyze temporal variations in communication performance, this sequence is divided into sliding windows of length $W$, shifted by $F$ steps. Given the total sequence length $l$, the number of windows $D$ is expressed as
$D=\frac{l+F-W}{F}.$ (2)
@@
The energy consumption model of a LoRa device used in this paper models the energy consumption during data communication in active mode, which can be calculated below [4].
@@ III-A Schwarz Information Criterion
A change is detected when the following condition is satisfied:
@@
where $\theta$ is an empirically determined detection threshold. When this condition holds at time $t$, the ED resets its learning history for all parameter combinations $k_{i}\in K$ in the UCB1-tuned algorithm, including cumulative reward $R_{k_{i}}(t)$, selection count $N_{k_{i}}(t)$, variance estimate $V_{k_{i}}(t)$, UCB score $P_{k_{i}}(t)$, and ACK sequence $\Omega_{k_{i}}$, the values are reset to zero when an environmental change is detected. This enables each ED to promptly adapt to new communication environments without being affected by learning histories.
@@ III-B SIC-based Transmission Parameter Selection
1:$t=0$, $P_{k_{i}}(t)=0$, $R_{k_{i}}(t)=0$, $N_{k_{i}}(t)=0$, $V_{k_{i}}(t)=0$ for all $k_{i}\in K$, $\Omega=\emptyset$
@@
4:  Append $s_{i}$ to history $\Omega$
@@
12:    $r\leftarrow\dfrac{Payload_{k^{*}}(t)}{E_{\mathrm{ToA}}}$
@@
16:  Update $R_{k^{*}}(t)$, $N_{k^{*}}(t)$, and $V_{k^{*}}(t)$ using reward $r$
17:  Compute $\mathrm{SIC}(D)$ and $\mathrm{SIC}(j)$ using  (12)–(13)
@@
20:    Reset history $\Omega\leftarrow\emptyset$
@@
The first term represents the empirical mean reward, and the second term promotes exploration of under-selected actions. The parameter combination with the largest $P_{k_{i}}(t)$ is selected for the next transmission (line 7). After each transmission (lines 8–14), the ED observes the ACK result $s_{t}\in\{0,1\}$ and measures the total active energy $E_{\mathrm{ToA}}$. If $s_{t}=1$, the reward is calculated as
$r=\frac{\text{Payload}_{k^{*}}(t)}{E_{\text{ToA}}},$ (16)
corresponding to a successful transmission; otherwise, $r=0$ for a failed one. The obtained reward is then used to update the cumulative reward $R_{k^{*}}(t)$, selection count $N_{k^{*}}(t)$, and variance estimate $V_{k^{*}}(t)$ (line 15). Next, SIC is applied to the ACK reception history $\Omega$ to detect potential environmental changes (lines 16–20). When the threshold condition defined in Eq. (14) is satisfied, the algorithm resets all learning variables and the ACK history to remove outdated information. The process continues until the maximum transmission count is reached. The subroutine UCB1-TUNED (lines 23–28) computes the UCB score for each $k_{i}$ using the current statistics and returns the combination $k^{*}$ with the highest score [14].

### v3

# Schwarz Information Criterion Aided MAB for Resource Allocation in Dynamic LoRa System

@@ I Introduction
With the rapid proliferation of the Internet of Things (IoT) devices in recent years, it has become increasingly common for a massive number of terminals to transmit data simultaneously [1]. This explosive growth has intensified channel contention in Low-Power Wide-Area Networks (LPWANs), particularly in Long Range (LoRa) networks, which stand out as a prominent example of LPWAN technology [2, 3]. To use limited wireless resources efficiently, each LoRa device need to optimize its transmission parameters in response to its dynamically changing communication environment.
@@
One representative decentralized technique used in distributed resource allocation is the Multi-Armed Bandit (MAB) algorithms [6, 7, 8, 9, 10, 11]. MAB is a reinforcement learning method that balances exploration and exploitation to learn the optimal choice, and it has been well applied to transmission parameter optimization. In particular, Upper Confidence Bound (UCB)1-tuned algorithm, which considers the variance of rewards for each option, has been widely used due to its enhanced exploration efficiency. Such decentralized online learning mechanisms can be regarded as a lightweight form of AI-driven networking, where EDs autonomously learn and adapt their transmission policies based on local observations. However, most MAB based distributed transmission parameter selection methods, including UCB1-tuned, rely heavily on historical observations. This dependence results in delayed adaptation when the communication environment changes abruptly, as these algorithms lack mechanisms to automatically discard past learning history [12]. Consequently, performance degradation in dynamic environments becomes unavoidable.
@@
  • We propose a lightweight distributed learning based transmission parameter selection method. In our proposed method, only a small set of statistics need to be stored, and requires neither complex matrix operations nor centralized control. Consequently, it incurs low computational and memory overhead, making it well suited for implementation on memory-constrained IoT devices.
@@ II System Model
The LoRa system considered in this study is illustrated in Fig. 1, consisting of a single GateWay (GW), multiple EDs, and an external interference source (jammer node). In the LoRa system, multiple channels are available in the network, and each ED selects one of them to transmit data to the GW. However, when the jammer occupies or blocks a specific channel, transmissions over that channel are more likely to experience interference. Therefore, we assume a dynamic communication environment where the jammer intermittently affects part of the frequency spectrum, causing certain channels to transition from an available state to an unavailable state during operation. Each ED needs to adaptively select transmission parameters that achieve the highest energy efficiency under varying channel conditions.
Let $G$ be the number of LoRa EDs, $M$ be the number of available channels, $P$ be the number of transmission power (TP) levels, and $H$ be the number of bandwidth (BW) options. The set of available channels is denoted as $C=\{c_{1},c_{2},\ldots,c_{m},\ldots,c_{M}\}$, the set of TP levels as $U=\{u_{1},u_{2},\ldots,u_{p},\ldots,u_{P}\}$, and the set of BW options as $B=\{b_{1},b_{2},\ldots,b_{h},\ldots,b_{H}\}$. Let $K=\{k_{1},k_{2},\ldots,k_{m\times p\times h},\ldots,k_{M\times P\times H}\}$ be the set of all possible combinations of channel, TP, and BW. Each ED selects one parameter set from $K$ before transmitting data based on our proposed method.
@@
Furthermore, each ED records the presence or absence of ACKs for each transmission and accumulates this as a transmission history. In this study, we introduce a mechanism that statistically detects changes in the communication environment based on this history using the SIC. When a change is detected, the past learning history is discarded, and the learning process is reinitialized to enable quick adaptation to the new environment. Each ED maintains $M$ binary observation sequences of ACK receptions, and the $m$-th sequence corresponds to $c_{m}$ is represented as follows:
$\Omega_{c_{m}}=\{\,s_{i}\mid s_{i}\in\{0,1\},\,i=1,2,\ldots,l\,\},$ (1)
where $l$ denotes the length of the $\Omega_{c_{m}}$, $s_{i}=1$ denotes a successful ACK reception, and $s_{i}=0$ denotes a failure. Let $l_{\max}$ denote the maximum length of each binary observation sequence. Historical data prior to $l_{\max}$ are discarded to maintain low memory overhead and real-time performance. To analyze temporal variations in communication performance, this sequence is divided into sliding windows of length $W$, shifted by $F$ steps. Given the total sequence length $l$, the number of windows $D$ is expressed as:
$D=\lfloor\frac{l+F-W}{F}\rfloor.$ (2)
@@
The energy consumption model of a LoRa ED used in this paper models the energy consumption during data communication in active mode, which can be calculated below [4].
@@ III-A Schwarz Information Criterion
An environmental change is detected when the following condition is satisfied:
@@
where $\theta$ is an empirically determined detection threshold. Smaller $\theta$ adapts faster but risks false resets; larger $\theta$ is steadier but slower to respond. When this condition holds at time $t$, the ED resets its learning history for all parameter combinations $k_{i}\in K$ in the UCB1-tuned algorithm, including cumulative reward $R_{k_{i}}(t)$, selection count $N_{k_{i}}(t)$, variance estimate $V_{k_{i}}(t)$, UCB score $P_{k_{i}}(t)$ to zero. In addition, the ACK sequences $\Omega_{c_{m}}$ for for all $c_{m}\in\mathcal{C}$ are reset to empty set. This enables each ED to promptly adapt to new communication environments without being affected by learning histories.
@@ III-B SIC-based Transmission Parameter Selection
1:$t=0$, $P_{k_{i}}(t)=0$, $R_{k_{i}}(t)=0$, $N_{k_{i}}(t)=0$, $V_{k_{i}}(t)=0$ for all $k_{i}\in K$, $\Omega_{c_{m}}\leftarrow\emptyset$ for all $c_{m}\in\mathcal{C}$
@@
4:  Append $s_{i}$ to the corresponding history $\Omega_{c_{m}}$
@@
12:    $\Delta R\leftarrow\dfrac{Payload_{k^{*}}(t)}{E_{\mathrm{ToA}}}$
@@
16:  Update $R_{k^{*}}(t)$, $N_{k^{*}}(t)$, and $V_{k^{*}}(t)$ based on the reward $\Delta R$
17:  Compute $\mathrm{SIC}(D)$ and $\mathrm{SIC}(j)$ using  (12)–(13) based on the $\Omega_{c_{m}}$ that corresponds to the selected channel $c_{m}$
@@
20:    Reset history $\Omega_{c_{m}}\leftarrow\emptyset$ for all $c_{m}\in\mathcal{C}$
@@
The first term represents the empirical mean reward, and the second term promotes exploration of under-selected actions. The parameter combination with the largest $P_{k_{i}}(t)$ is selected for the next transmission (line 7). After each transmission (lines 8–14), the ED observes the ACK result and calculates reward. If the ACK result $s_{t}=1$, the reward is calculated as
$\Delta R=\frac{\text{Payload}_{k^{*}}(t)}{E_{\text{ToA}}},$ (16)
corresponding to a successful transmission; otherwise, $\Delta R=0$ for a failed one. Then, $R_{k^{*}}(t)$, $N_{k^{*}}(t)$, and $V_{k^{*}}(t)$ are updated based on the obtained reward and the observed ACK information (line 15). Next, SIC is applied to the ACK reception history $\Omega_{c_{m}}$ to detect potential environmental changes (lines 16–20). When the threshold condition defined in Eq. (14) is satisfied, the algorithm resets all learning variables and the ACK history to remove outdated information. The process continues until the maximum transmission count is reached. The subroutine UCB1-TUNED (lines 23–28) computes the UCB score for each $k_{i}$ using the current statistics and returns the combination $k^{*}$ with the highest score [14].